# Recursive Joint Cross-Modal Attention for Multimodal Fusion in Dimensional Emotion Recognition

### v2

@@ 4.2.2 Training Details
For visual modality, random flipping, and random crop with a size of $40$ are used for data augmentation in training, while only center crop is used for validation. For audio and visual features, the input data is normalized in order to have a mean and standard deviation of $0.5$. For text modality, the BERT features are normalized to ensure the mean as 00 and standard deviation as $1$. Adam optimizer is used with a weight decay of $0.001$ and batch size is set to be $12$. The models are trained separately for valence and arousal. The maximum number of epochs is set to be $100$ and early stopping is employed to avoid over-fitting. The hyperparameters of initial learning rate and minimum learning rate are set to be $51e-$ 5 and $81e-$ 8 respectively. In our training strategy, we have deployed warm up scheme using $ReduceLROnPlateau$ scheduler with a patience of $5$ and factor of $0.1$ based on the CCC score of validation partition. It has been shown that gradual training of backbones of individual modalities along with the fusion model by gradually fine-tuning the layers of the backbones helps to improve the performance of the system [41]. Therefore, we have deployed a similar strategy in our training framework, where three groups of layers for visual (Resnet-50) and audio (VGG) backbones are progressively selected for fine-tuning. Initially at epoch 00, the first group is unfrozen and the learning rate is linearly warmed up to $51e-$ 5 within an epoch. Then repetitive warm-up is employed until epoch $5$, after which $ReduceLROnPlateau$ is used to update the learning rate. The learning rate is gradually dropped with a factor of $0.1$ until validation CCC does not improve over 5 consecutive epochs. After which the second group is unfrozen and the learning rate is reset to $51e-$ 5, followed by warm-up scheme with $ReduceLROnPlateau$. The procedure is repeated till all the layers are fine-tuned for audio and visual backbones. Also, note that best model state dictionary over prior epochs is loaded at the end of each epoch to mitigate the issues of over-fitting. To further control the problem of over-fitting, we have employed cross-validation with 6 folds, where fold 00 partition is the same as the original partition provided by the organizers [59]. The results obtained from the 6-fold cross-validation are shown in Table 1. In all these experiments, we have used 3 iterations in the fusion model (i.e., $l$=3).
Table 1: CCC of the proposed approach on the six folds of cross-validation. Bold indicates highest scores. Fold 0 is the official validation set.
@@ 4.2.3 Loss Function
The concordance correlation coefficient ($\rho_{c}$) is the widely-used evaluation metric in the literature for DER to measure the level of agreement between the predictions ($x$) and ground truth ($y$) annotations [11]. Let $\mu_{x}$, and $\mu_{y}$ represent the mean of predictions and ground truth, respectively. Similarly $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ denote the variance of predictions and ground truth, respectively, then $\rho_{c}$ between the predictions and ground truth can be obtained as:
@@
where $\sigma_{xy}^{2}$ denotes the covariance between predictions and ground truth. Though Mean Square Error (MSE) is the commonly used loss function for regression models, we use CCC-based loss function as it is a standard loss function in the literature for DER [9, 11], which is given by
@@
Table 3: CCC of the proposed fusion model by varying the number of recursions. Bold indicates highest scores.
@@ 5 Results and Discussion
Method Modalities Valence Arousal Mean
Netease Fuxi AI Lab [61] Audio, Visual 0.6873 0.6569 0.6721
RJCMA (Ours) Audio, Visual, Text 0.5418 0.6196 0.5807
CtyunAI [62] Visual 0.5223 0.6057 0.5640
SUN-CE [63] Audio, Visual 0.5355 0.5861 0.5608
USTC-IAT-United [64] Audio, Visual 0.5208 0.5748 0.5478
HSEmotion [65] Visual 0.4925 0.5461 0.5193
KBS-DGU Audio, Visual 0.4836 0.5318 0.5077
ETS-LIVIA [66] Audio, Visual 0.4198 0.4669 0.4434
CAS-MAIS Audio, Visual, Text 0.4245 0.3414 0.3830
IMLAB [67] Visual 0.2912 0.2456 00.2684
Baseline [59] Visual 0.2110 0.1910 0.2010
@@ 5.1 Ablation Study
In order to understand the impact of recursive mechanism, we have conducted a series of experiments on the validation set by varying the number of recursions as shown in Table 3. First, we did an experiment with a single recursion, which is the same as joint cross attention [9] for audio, visual and text modalities. Now we executed multiple experiments by slowly increasing the number of recursions and found that the performance of the system gradually increases with multiple recursions. This shows that recursive mechanism helps in obtaining more robust feature representations by progressively refining the features. We have achieved best results at 3 iterations, beyond that the performance of the system declines. We hypothesize that this can be attributed to the fact that though recursive fusion initially helps to improve the performance, it may result in over-fitting with more iterations, resulting in a performance decline on the validation set. A similar trend of performance improvement with multiple recursions is also observed for other audio-visual tasks such as person verification [25] and event localization [24], thereby underscoring our hypothesis.
@@ 5.2 Comparison to State-of-the-art
Most of the approaches evaluated on the Affwild2 dataset have been submitted to previous ABAW challenges. Therefore, we have compared the performance of the proposed approach with some of the relevant state-of-the-art models of previous ABAW challenges as shown in Table 2. Several approaches explored ensemble-based methods by using multiple encoders for each modality, followed by transformers to encode the concatenated multimodal feature representation [36, 18, 17]. By exploiting multiple backbones and large-scale training with external datasets, [17] significantly improved the test set performance on both valence and arousal. Similarly, [36] also explored multiple backbones and showed a better improvement in the arousal performance of the test set than [17]. Even though ensemble-based methods show better performance on test set, it is often cumbersome and computationally expensive. Zhang et al. [37] showed that exploring Masked Auto-Encoders (MAE) is a promising line of research to achieve better generalization and consistently improved performance in both valence and arousal. Zhang et al. [47] explored the leader-follower attention model for multimodal fusion, where audio and text modalities are leveraged to attend to the visual modality, and showed good performance on the test set without the need for multiple backbones for each modality.
Praveen et al. [9] proposed Joint Cross Attention (JCA) by introducing joint feature representation in the cross-attention framework and showed significant improvement in the validation set, especially for valence. They further improved the performance of their approach by deploying recursive mechanism and LSTMs for temporal modeling of individual modalities and multimodal feature representations [23]. However, they do not seem to have a better generalization ability as they fail to show improvement in the test set. We hypothesize that this may be due to naive audio and visual backbones. Therefore, to have a fair comparison with [23], we have reimplemented their fusion model by replacing their audio and visual backbones with Vggish [43] and Resnet-50 pre-trained on MSCELEB-1M and FER+ [41], followed by TCNs. By replacing the visual and audio backbones of [23] and progressive fine-tuning of the backbones similar to that of [41], the problem of over-fitting has been mitigated and showed improvement in test set performance also. We can observe that performance has been improved in both validation and test sets by introducing a text modality into the RJCA framework work [23], especially for arousal. Note that even though the performance of valence for [37, 47] as well as our approach on the official validation set is lower, they achieved better performance on other folds of cross-validation, thereby improving the test set performance.
@@ 5.3 Results on Test Set
Multimodal learning is gaining a lot of attention for affective behavior analysis as most of the methods submitted for the valence-arousal challenge of 6th ABAW competition [59] employed multimodal fusion [61, 63, 64, 66]. Several methods explored ensemble-based fusion for better generalization ability [66, 61, 63]. Another widely-explored strategy to improve the test set performance is to exploit pre-training with multiple large-scale datasets using Masked Auto-Encoders (MAE) [61, 62] or multiple backbones for each modality [63, 65, 64]. Netease Fuxi AI Lab [61] used MAEs pre-trained with 5 external datsets of around 262M images and Vggish model [43] for audio modality, followed by ensemble-based fusion and showed significant improvement for both valence and arousal. Similarly, CtyunAI [62] also used MAE for visual modality pre-trained with 4 external datasets, followed by TCNs for temporal modeling and achieved very good performance for arousal than valence. Sun-CE [63] explored multiple backbones for both audio and visual modalities, followed by ensemble-based fusion, while USTC-IAT-United [64] used multiple backbones only for audio modality and achieved decent performance for both valence and arousal. KBS-DGU and ETS-LIVIA [66] exploited ensemble of self-attention and cross-attention models for improving the performance of audio-visual fusion. Unlike other approaches, we have explored the prospect of simultaneously capturing the intra- and inter-modal relationships across the modalities in a recursive fashion and achieved very good performance for both valence and arousal without using multiple backbones or pre-training with large-scale external datasets. Therefore, the performance of our approach can be solely attributed to the robustness of the sophisticated fusion model, providing a cost-effective solution for affective behavior analysis in-the-wild.
@@ 6 Conclusion
In this work, we showed that effectively capturing the synergic relationships pertinent to the intra- and inter-modal characteristics among the audio, visual and text modalities can significantly improve the performance of the system. By introducing the joint representation in the cross-attentional framework, we are able to simultaneously capture both intra- and inter-modal relationships across the audio, visual and text modalities. The performance of the system is further enhanced using recursive fusion by progressively refining the features to obtain robust multimodal feature representations. Experimental results on the challenging Affwild2 dataset indicate that the proposed model is able to achieve superior performance of multimodal fusion, outperforming most of the methods for both valence and arousal. The performance of the system can also be improved by leveraging advanced text encoders and sophisticated backbones for the individual modalities.

### v4

@@ 4.2.2 Training Details
For visual modality, random flipping, and random crop with a size of $40$ are used for data augmentation in training, while only center crop is used for validation. For audio and visual features, the input data is normalized in order to have a mean and standard deviation of $0.5$. For text modality, the BERT features are normalized to ensure the mean as 00 and standard deviation as $1$. Adam optimizer is used with a weight decay of $0.001$ and batch size is set to be $12$. The models are trained separately for valence and arousal. The maximum number of epochs is set to be $100$ and early stopping is employed to avoid over-fitting. The hyperparameters of the initial learning rate and minimum learning rate are set to be $51e-$ 5 and $81e-$ 8 respectively. In our training strategy, we have deployed a warm-up scheme using $ReduceLROnPlateau$ scheduler with patience of $5$ and a factor of $0.1$ based on the CCC score of the validation partition. It has been shown that gradual training of the backbones of individual modalities along with the fusion model by gradually fine-tuning the layers of the backbones helps to improve the performance of the system [42]. Therefore, we have deployed a similar strategy in our training framework, where three groups of layers for visual (Resnet-50) and audio (VGG) backbones are progressively selected for fine-tuning. Initially at epoch 00, the first group is unfrozen and the learning rate is linearly warmed up to $51e-$ 5 within an epoch. Then repetitive warm-up is employed until epoch $5$, after which $ReduceLROnPlateau$ is used to update the learning rate. The learning rate is gradually dropped with a factor of $0.1$ until validation CCC does not improve over 5 consecutive epochs. After which the second group is unfrozen and the learning rate is reset to $51e-$ 5, followed by a warm-up scheme with $ReduceLROnPlateau$. The procedure is repeated till all the layers are fine-tuned for audio and visual backbones. Also, note that the best model state dictionary over prior epochs is loaded at the end of each epoch to mitigate the issues of over-fitting. To further control the problem of over-fitting, we have employed cross-validation with 6 folds, where the fold 00 partition is the same as the original partition provided by the organizers [60]. The results obtained from the 6-fold cross-validation are shown in Table 1. In all these experiments, we have used 3 iterations in the fusion model (i.e., $l$=3).
Table 1: CCC of the proposed approach on the six folds of cross-validation. Bold indicates the highest scores. Fold 0 is the official validation set.
@@ 4.2.3 Loss Function
The concordance correlation coefficient ($\rho_{c}$) is the widely-used evaluation metric in the literature for DER to measure the level of agreement between the predictions ($x$) and ground truth ($y$) annotations [10]. Let $\mu_{x}$, and $\mu_{y}$ represent the mean of predictions and ground truth, respectively. Similarly, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ denote the variance of predictions and ground truth, respectively, then $\rho_{c}$ between the predictions and ground truth can be obtained as:
@@
where $\sigma_{xy}^{2}$ denotes the covariance between predictions and ground truth. Though Mean Square Error (MSE) is the commonly used loss function for regression models, we use the CCC-based loss function as it is a standard loss function in the literature for DER [21, 10], which is given by
@@
Table 3: CCC of the proposed fusion model by varying the number of recursions. Bold indicates the highest scores. Fold 1 is used for experiments with multiple recursions.
@@ 5 Results and Discussion
Method Modalities Ensemble Valence Arousal Mean
Netease Fuxi AI Lab [62] Audio, Visual Yes 0.6873 0.6569 0.6721
RJCMA (Ours) Audio, Visual, Text No 0.5418 0.6196 0.5807
CtyunAI [63] Visual No 0.5223 0.6057 0.5640
SUN-CE [64] Audio, Visual Yes 0.5355 0.5861 0.5608
USTC-IAT-United [65] Audio, Visual No 0.5208 0.5748 0.5478
HSEmotion [66] Visual No 0.4925 0.5461 0.5193
KBS-DGU Audio, Visual No 0.4836 0.5318 0.5077
ETS-LIVIA [67] Audio, Visual No 0.4198 0.4669 0.4434
CAS-MAIS Audio, Visual, Text No 0.4245 0.3414 0.3830
IMLAB [68] Visual No 0.2912 0.2456 00.2684
Baseline [60] Visual No 0.2110 0.1910 0.2010
@@ 5.1 Ablation Study
In order to understand the impact of the recursive mechanism, we have conducted a series of experiments by varying the number of recursions as shown in Table 3. First, we did an experiment with a single recursion, which is the same as joint cross attention [21] for audio, visual, and text modalities. Now we executed multiple experiments by slowly increasing the number of recursions and found that the performance of the system gradually increases with multiple recursions. This shows that the recursive mechanism helps in obtaining more robust feature representations by progressively refining the features. We have achieved the best results at 3 iterations, beyond that the performance of the system declines. We hypothesize that this can be attributed to the fact that though recursive fusion initially helps to improve the performance, it may result in over-fitting with more iterations, resulting in a performance decline on the validation set. A similar trend of performance improvement with multiple recursions is also observed for other audio-visual tasks such as person verification [25] and event localization [24], thereby underscoring our hypothesis.
@@ 5.2 Comparison to State-of-the-art
Most of the approaches evaluated on the Affwild2 dataset have been submitted to previous ABAW challenges. Therefore, we have compared the performance of the proposed approach with some of the relevant state-of-the-art models of previous ABAW challenges as shown in Table 2. Several approaches explored ensemble-based methods using multiple encoders for each modality, followed by transformers to encode the concatenated multimodal feature representation [36, 17, 16]. By exploiting multiple backbones and large-scale training with external datasets, Meng et al. [16] significantly improved the test set performance on both valence and arousal. Similarly, Zhou et al. [36] also explored multiple backbones and showed better improvement in the arousal performance of the test set than [16]. Even though ensemble-based methods show better performance on test set, it is often cumbersome and computationally expensive. Zhang et al. [37] showed that exploring Masked Auto-Encoders (MAE) is a promising line of research to achieve better generalization and consistently improved performance in both valence and arousal. Zhang et al. [42] explored the leader-follower attention model for multimodal fusion, where audio and text modalities are leveraged to attend to the visual modality, and showed good performance on the test set without the need for multiple backbones for each modality.
Praveen et al. [21] proposed Joint Cross Attention (JCA) by introducing a joint feature representation in the cross-attention framework and showed significant improvement in the validation set, especially for valence. They further improved the performance of their approach by deploying a recursive mechanism and LSTMs for the temporal modeling of individual modalities and multimodal feature representations [23]. However, they do not seem to have a better generalization ability as they fail to show improvement in the test set. We hypothesize that this may be due to naive audio and visual backbones, fine-tuned on Affwild2 dataset. Therefore, to have a fair comparison with [23], we have reimplemented their fusion model by replacing their audio and visual backbones with Vggish [44] and Resnet-50 pre-trained on MSCELEB-1M and FER+ [42], followed by TCNs. By replacing the visual and audio backbones of [23] and progressive fine-tuning of the backbones similar to that of [42], the problem of over-fitting has been mitigated and showed improvement in test set performance also. We can observe that performance has been improved in both validation and test sets by introducing text modality into the RJCA framework [23], especially for arousal. Note that even though the performance of the valence for [37, 42] as well as our approach on the official validation set is lower, they achieved better performance on other folds of cross-validation, thus improving the performance of the test set.
@@ 5.3 Results on Test Set
Multimodal learning is gaining a lot of attention for affective behavior analysis as most of the methods submitted for the valence-arousal challenge of the 6th ABAW competition [60] employed multimodal fusion [62, 64, 65, 67]. Some of these methods explored ensemble-based fusion for better generalization ability [62, 64]. Another widely explored strategy to improve the test set performance is to exploit pre-training with multiple large-scale datasets using Masked Auto-Encoders (MAE) [62, 63] or multiple backbones for each modality [64, 66, 65]. Netease Fuxi AI Lab [62] used MAEs pretrained with 5 external data sets of around 262M images and Vggish model [44] for the audio modality, followed by ensemble-based fusion, and showed a significant improvement for both valence and arousal, achieving first place in the challenge. Similarly, CtyunAI [63] also used MAE for visual modality pre-trained with 4 external datasets, followed by TCNs for temporal modeling, and achieved very good performance for arousal than valence. Sun-CE [64] explored multiple backbones for both audio and visual modalities, followed by ensemble-based fusion, while USTC-IAT-United [65] used multiple backbones only for the audio modality and achieved decent performance for both valence and arousal. KBS-DGU and ETS-LIVIA [67] exploited combining features from multiple fusion models of self-attention and cross-attention models for improving the performance of audio-visual fusion. KBS-DGU explored both self-attention and cross-attention models, while ETS-LIVIA [67] used joint feature representation with cross-attention models across audio and visual modalities. Unlike other approaches, we have explored the prospect of simultaneously capturing the intra- and inter-modal relationships across audio, visual, and text modalities recursively and achieved very good performance for both valence and arousal without using multiple backbones or pre-training with large-scale external datasets. Therefore, the performance of our approach can be solely attributed to the robustness of the sophisticated fusion model, providing a cost-effective solution for affective behavior analysis in-the-wild.
@@ 6 Conclusion
In this work, we showed that effectively capturing the synergic relationships pertinent to the intra- and inter-modal characteristics among the audio, visual, and text modalities can significantly improve the performance of the system. By introducing the joint representation in the cross-attentional framework, we can simultaneously capture both intra- and inter-modal relationships across the audio, visual, and text modalities. The performance of the system is further enhanced using recursive fusion by progressively refining the features to obtain robust multimodal feature representations. Experimental results in the challenging Affwild2 dataset indicate that the proposed model can achieve a better multimodal fusion performance, outperforming most methods for both valence and arousal. The performance of the system can also be improved by leveraging advanced text encoders and sophisticated backbones for the individual modalities.
Acknowledgements The authors wish to acknowledge the funding from the Government of Canada’s New Frontiers in Research Fund (NFRF) through grant NFRFR-2021-00338 and Ministry of Economy and Innovation (MEI) of the Government of Quebec for the continued support.